xx CalcK — Calculate K_{ij}

CalcK — Calculate

Erik Schnetter <schnetter@aei.mpg.de>

Date

Abstract

Calculate the extrinsic curvature from several time levels of the three-metric , lapse , and shift .

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11calcxextcurv 10Scope: privatex01 BOOLEAN

11Description: Calculate extrinsic curvature

11                     10  x01  Default: no

11

11copyxtoxnext  10Scope: privatex01  BOOLEAN

11Description: Copy ADMBase::metric to CalcK::metricxnext

11                     10  x01  Default: no

11

11copyxtoxprev  10Scope: privatex01  BOOLEAN

11Description: Copy ADMBase::metric to CalcK::metricxprev

11                     10  x01  Default: no

11

11copyxtoxprev2  10Scope: privatex01  BOOLEAN

11Description: Copy ADMBase::metric to CalcK::metricxprev2

11                     10  x01  Default: no

11

11extcurvxboundary  10Scope: privatex01  STRING

11Description: Boundary condition for the extrinsic curvature

11 Range 10 x01 Default: Robin
11.*  10must be a registered boundary condition

11

11extcurvxboundaryxoptions 10Scope: privatex01  STRING

11Description: Options for the boundry condition for the extrinsic curvature

11 Range 10 x01 Default: decayxpower=2.0
11.*  10must have the options table syntax

11

11metricxtype  10Scope: shared from ADMBASEx01KEYWORD

11

7 Interfaces

General Implements: calck

Inherits: admbase staticconformal Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 metricxprev 10  x01  compact110
11                 10  gxxxprev  x01 description11ADM 3-metric
11                 10  x  x01 description11previous time
11                 10  gxyxprev  x01 dimensions113
11                 10  gxzxprev  x01 distribution11DEFAULT
11                 10  gyyxprev  x01 group type11GF
11                 10  gyzxprev  x01 timelevels111
11                 10  gzzxprev  x01variable type11REAL
11 metricxprev2 10  x01  compact110
11                 10  gxxxprev2  x01 description11ADM 3-metric
11                 10  x  x01 description11pre-previous time
11                 10  gxyxprev2  x01 dimensions113
11                 10  gxzxprev2  x01 distribution11DEFAULT
11                 10  gyyxprev2  x01 group type11GF
11                 10  gyzxprev2  x01 timelevels111
11                 10  gzzxprev2  x01variable type11REAL
11 metricxnext 10  x01  compact110
11                 10  gxxxnext  x01 description11ADM 3-metric
11                 10  x  x01 description11next time
11                 10  gxyxnext  x01 dimensions113
11                 10  gxzxnext  x01 distribution11DEFAULT
11                 10  gyyxnext  x01 group type11GF
11                 10  gyzxnext  x01 timelevels111
11                 10  gzzxnext  x01variable type11REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn EinsteinAnalysis/CalcK. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage

11Always:  11x
11metricxprev metricxprev2 metricxnext11x

11x  11x

11

Scheduled Functions

CCTKxPOSTINITIAL (conditional) x calckxcopyxtoxprev x copy admbase::metric to calck::metricxprev x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxPOSTINITIAL (conditional) x calckxcopyxtoxprev2 x copy admbase::metric to calck::metricxprev2 x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxPOSTINITIAL (conditional) x calckxcopyxtoxnext x copy admbase::metric to calck::metricxnext x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxPOSTINITIAL (conditional) x calck x calculate kxij from gammaxij, alpha, and betax x

11x11Language:11fortran
11 x 11Sync:  11admbase::curv
11 x 11Type: 11function

11

CCTKxPOSTINITIAL (conditional) x applybcs x apply boundary conditions x

11x11After:11calck
11 x 11Type:11group

Aliased Functions x

11Alias Name: xxxxxxx11Function Name:
11ApplyBCs  11CalcKxApplyBCs